Please assume that heat is added to a moveable piston which has a constant pressure of 200 ?pa. Assuming that the volume of gas contained in this cylinder expands from 0.1 ?3 to 0.2 ?3, please calculate the amount of work done by the gas contained within this piston.
Answer #1

Work done by gas for closed system is (pdv) where dv is change in volume

Now for constant pressure process work done =p dv

= p*(vf-vi)

where vf and vi are final and initial volumes

So W = 200*103*(0.2-0.1)= 20000 J or 20KJ

Work is positive means it is done by the system